hypotenuse is 2.65m in a right angle triangle. Neither leg is given a length but we know that the legs are the same length. How would I solve this?
Answer by CharlesG2(834)

hypotenuse is 2.65m in a right angle triangle. Neither leg is given a length but we know that the legs are the same length. How would I solve this?

legs equal --> leg a = leg b --> a = b
hypotenuse = c = 2.65 m
Pythagorean Theorem --> a^2 + b^2 = c^2 --> 2a^2 = c^2, since a = b

2a^2 = 2.65^2
2a^2 = 7.0225
a^2 = 3.51125
a = b = 1.873833 to 6 places
a = b = 1.87 m to 2 places